OpenStudy (anonymous):

2x=5y=20,3x-10y=37 elimination

OpenStudy (anonymous):

i keep getting confused

OpenStudy (anonymous):

is that 2x-5y = 20?

OpenStudy (anonymous):

+5y

OpenStudy (anonymous):

2x+5y=20 3x-10y=37 I personally recommend getting rid of the y because you just have to multiply top equation by 2 4x+10y=40 3x-10y=37 This time you add to cancel out the y because +10 + (-10) = 0 7x=77 x=11

OpenStudy (anonymous):

Do you get it? :o

OpenStudy (anonymous):

alot better now that you wannt step by steo thanks again lol